# Bibliografia.bib

### v1

@@ Abstract
Blockchain technology has revolutionized financial markets by enabling decentralized exchanges (DEXs) that operate without intermediaries. Uniswap V2, a leading DEX, facilitates the rapid creation and trading of new tokens, offering high return potential but exposing investors to significant risks. In this work, we analyze the financial impact of newly created tokens, assessing their market dynamics, profitability and liquidity manipulations. Our findings reveal that a significant portion of market liquidity is trapped in honeypots, reducing market efficiency and misleading investors. Applying a simple buy-and-hold strategy, we are able to uncover some major risks associated with investing in newly created tokens, including the widespread presence of rug pulls and sandwich attacks. We extract the optimal sandwich amount, revealing that their proliferation in new tokens stems from higher profitability in low-liquidity pools. Furthermore, we analyze the fundamental differences between token price evolution in swap time and physical time. Using clustering techniques, we highlight these differences and identify typical patterns of honeypot and sellable tokens. Our study provides insights into the risks and financial dynamics of decentralized markets and their challenges for investors.
@@ 1 Introduction
The advent of blockchain technology has revolutionized financial markets, enabling decentralized platforms that operate independently of centralized intermediaries. Among these innovations, Decentralized Exchanges (DEXs) have emerged as key players in the cryptocurrency ecosystem. Unlike centralized exchanges, DEXs rely on decentralized protocols implemented via smart contracts deployed on blockchains, allowing trustless, peer-to-peer trading without the need for custodial services. This framework has led to the proliferation of cryptocurrency tokens and trading strategies, driven by the innovative mechanism of Automated Market Makers (AMMs)[AMMeco], which replaces the traditional order book model with a liquidity pool-based system. One notable example is Uniswap V2 [uni2core, analysisUniswap], a widely-used DEX protocol that supports the creation of liquidity pools for token pairs, typically involving Ethereum (ETH). On average, $\sim 15$ new tokens paired with Ethereum are introduced hourly on Uniswap V2 between October 10, 2024, and December 2, 2024. Many of these tokens initially exhibit extremely high returns, creating an illusion of huge profits and attracting inexperienced investors. However, beneath this apparent simplicity and profitability lie significant risks and vulnerabilities that can lead to partial or total loss of funds. Indeed, the decentralized nature of DEXs introduces a set of security challenges, which have been extensively analyzed in the context of computer security and formal verification methods [Zhou2022SoKDF]. These studies highlight that these coins and many investments hide important vulnerabilities, exposing users to risks that require careful evaluation.
@@
Contribution of the work. To our knowledge, most existing studies on the vulnerabilities and risks associated with decentralized exchanges, including phenomena such as honeypots and backdoors, have primarily approached the topic from the perspective of computer security and formal verification. Although these works have significantly advanced methods to detect malicious contracts and identify vulnerabilities in smart contracts, they often lack a quantitative and financial perspective. Specifically, numerous methodologies have been proposed to identify honeypots or detect vulnerabilities, leveraging static and dynamic analysis techniques, formal methods, and auxiliary service data [Qin2021QuantifyingBE, Zhou2020HighFrequencyTO, HoneypotWeb1, HoneypotWeb2, TorresHoney, honeypot_url, Sun2024SoKCA]. However, there has been little to no contribution focused on understanding the broader financial impact of these phenomena, particularly in relation to the scale and significance of the proliferation of new tokens created within the decentralized finance ecosystem. In this work, we aim to bridge this gap by performing a quantitative and financial analysis of the creation and proliferation of new tokens on Uniswap V2.
@@ 2 Dataset and Token Classification
To conduct our analysis, we collected on-chain data from the Ethereum blockchain, focusing on newly created tokens and their trading activity on DEX. The dataset consists of 17,194 tokens launched between October 10, 2024, and December 2, 2024. Data collection was carried out by connecting to an Ethereum node using an RPC provider such as Infura or Alchemy via Web3. Newly created pairs were identified by monitoring blockchain events and tracking the creation of token pairs, recording their contract addresses along with the associated tokens (token_x, token_y). This allowed us to follow the introduction of new tokens into the ecosystem. We then collected corresponding swap events to capture the amounts swapped in and out, providing insights into price dynamics and trading activity. Additionally, we extracted liquidity-related data by examining mint and burn events, recording the amounts of liquidity tokens minted and burned to assess liquidity provisioning over time. This dataset serves as the foundation for analyzing token behavior, price evolution, and liquidity trends of early-stage token dynamics in DEX.
@@
We performed a time-based analysis, observing each token’s entire trading history to detect any transition into honeypot behavior. We leveraged external security detection services such as Honeypot.is, which specializes in identifying honeypots, and GoPlus Labs, which provides security scoring for tokens. These services allowed us to systematically verify whether a token exhibited selling restrictions at any point in its lifecycle. Based on this analysis, we categorized tokens at the end of their observed lifetime. Thus, we labeled as honeypots those coins that were either initially classified as such or became honeypots later in their lifecycle. However, our analysis indicates that the majority of honeypots transition into this state within the first few swaps. It is important to note that GoPlus and Honeypot.is use different methods to identify honeypots. The former analyzes the smart contract directly, while the latter simulates sell transactions to determine whether a token can be liquidated. In our study, we classify a token as a honeypot if at least one of these two providers flags it as such.
Tokens that never displayed honeypot behavior were classified as sellable, comprising approximately 2,080 tokens ($\simeq 12\%$ of the total). Conversely, tokens that exhibited locked or blocked selling at any stage were classified as honeypots, accounting for around 15,114 tokens ($\simeq 88\%$ of the total).
@@ 3 Early-Stage Properties of Newly Created Tokens
In this section, we analyze the early characteristics of newly created tokens on Uniswap V2, including their creation frequency, classification as sellable tokens or honeypots, and average lifespan. Honeypots significantly distort market dynamics by inflating trading volumes while restricting liquidity. By preventing selling, they create an illusion of high demand and price appreciation, misleading traders and absorbing substantial market volume. Understanding how much liquidity remains locked in these contracts provides insights into the true market potential of sellable tokens and the challenges faced by legitimate projects. A new token lifecycle begins with its creation on the Ethereum blockchain, followed by the establishment of a liquidity pool on Uniswap V2, where the owner pairs the token with ETH to enable trading. Initial trades occur almost immediately, often within the first block, driven by automated bots and, in some cases, the creator simulating demand. This early activity frequently causes a sharp price surge due to low liquidity. However, this surge is often short-lived, as subsequent selling pressure, sandwich attacks, and liquidity rug-pulls emerge. Analyzing these early-stage price dynamics, which is precisely what we do in this paper, helps distinguish between tokens with genuine growth potential and those exhibiting market manipulation.
@@
(b) Distribution of the number of new tokens created per hour.
Figure 1:
@@
(b) ETH amount distribution.
Figure 2:
@@
Figure 1(a) presents a histogram displaying the number of tokens created each day. Remarkably, the number of tokens created daily is on average around 200, highlighting the sustained activity in token creation. This trend underscores the significance of Uniswap V2 as a platform for launching and experimenting with new tokens. The orange part of the histogram represents the portion of honeypots among these newly created coins. The presence of a substantial fraction of honeypots within this continuous flow of token creation suggests that a significant portion of new projects are designed with malicious intent, aiming to trap liquidity rather than fostering genuine market participation. This dynamic not only skews the perception of overall market activity but also presents a major challenge for traders and investors, who must navigate an environment where a large share of tokens are inherently unsellable. As a direct consequence, a considerable portion of the liquidity circulating in the market is effectively locked within honeypot contracts, rendering it inaccessible to traders. This reduces the actual capital available for legitimate trading and investment, further distorting market efficiency. Thus, while the high daily token creation rate may signal strong market engagement, it also reflects an ongoing risk where a substantial amount of liquidity is continuously funneled into fraudulent contracts rather than supporting genuine projects. Figure 1(b) illustrates the distribution of the number of tokens created within one-hour bins. The peak of the distribution occurs around 15. However, the overall distribution is largely dominated by honeypots, which are created more frequently. In contrast, sellable tokens are launched at a significantly lower rate.
@@
A key question is how many swaps newly created tokens experience on average and whether they sustain trading activity or quickly become inactive. This behavior differs between honeypots and sellable tokens. Since honeypots prevent selling, they often see initial high activity as traders unknowingly buy in before realizing they cannot exit. In contrast, sellable tokens allow continuous trading, potentially leading to longer lifespans. Figure 2(a) confirms this: honeypots exhibit a sharp peak at low swap counts before rapidly decaying, indicating short-lived activity. sellable tokens, while also peaking at low swaps, show a broader distribution, suggesting prolonged trading. This is evident in the median swap count, which is 42 for honeypots but significantly higher at 363 for sellable tokens. Figure 2(b) presents the distribution in swap size in ETH. The sell (negative) side appears uniform, while the buy (positive) side shows distinct peaks at specific ETH values, likely due to bots executing trades with fixed investment amounts. This highlights that a substantial portion of market liquidity remains trapped in honeypots, misleading traders and distorting the perceived trading volume.
@@ 4 A buy-and-hold strategy on new tokens
In this section, we investigate the potential profitability of newly created tokens under the assumption that an investor is able to enter a position early in the token’s lifecycle—specifically, at the 60th swap. By systematically applying this strategy across the entire set of newly launched tokens, we can extract valuable insights into their financial behavior.
@@
Our trading bot operates in swap time. In particular, it observes transactions up to the 60th swap before performing its first purchase of tokens using an initial investment of $\Delta x=0.016$ ETH, buying an amount $\Delta y$ of the new token computed using the constant product rule formula
@@
where $x_{60}$ is the reserve of the newly created tokens in the liquidity pool at the 60th transaction, $y_{60}$ is the correspondent reserve of ETH in the liquidity pool and $r=0.997$ is the effective fee factor, accounting for the 0.3% swap fee.
After buying $\Delta x$ tokens, the strategy evaluates subsequent swap steps $n$ to determine an optimal time to sell. For each transaction $n>60$, the potential ETH receivable from selling the tokens ($\Delta x_{n}$) is calculated as
@@
where $x_{n}$ and $y_{n}$ are respectively the token reserves of the token and ETH in the pool at transaction $n>60$. The revenue $S_{n}$ at each transaction from the sale is computed as
@@
where $G$ is the gas fee incurred for each transaction that we take fixed to $G=0.002$ ETH. The strategy uses predefined thresholds to decide the swap time $n$ when to sell the tokens:
@@
The same is accounted if the order flow of the token ends before one of these conditions is reached.
Figure 3: Distribution of profits in ETH generated by our strategy. Among the 7,213 tokens in which the algorithm invested, 2,101 ($\sim 29.1\%$) failed to exit their position, resulting in a fixed loss of $P_{\text{loss}}=-0.018$ ETH. All remaining tokens yielded a positive profit. To improve visualization, we restrict the distribution to the range between 0 and 0.015 ETH. Beyond this threshold, there are an additional 162 tokens reaching a maximum profit of 1,056.07 ETH.
We apply this strategy to all the coins. The overall combined results indicate that $\sim$ 70.9% of the coins give a positive profit, while approximately $29.1\%$ of the tokens result in the fixed negative profit of $P_{\text{loss}}=-0.018$ ETH. The profit distribution for all tokens is shown in Figure 3, with the range restricted to $P\in[0,0.015]$ to improve visualization. We do not display the negative histogram at $P=P_{\text{loss}}$ and positive outliers. The maximum profit recorded on one coin is 1056.07 ETH. The total profit over all the coins, namely
@@
reaches 2180.24 ETH, with a total investment of 140.06 ETH, leading to an impressive percentage profit of 1556.67%. As expected, however, a large percentage of it is eaten by honeypots.
Indeed, focusing specifically on the honeypot results, we observe an even higher positive percentage of 71.83%, which includes the maximum profit of 1056.07 ETH. The total profit reaches 2163.81 ETH on a total investment of 104.88 ETH, leading to a percentage profit of 2063.13%. This clearly demonstrates that the majority of the apparent profit is generated by honeypot coins, creating the illusion of high profitability. However, this is entirely deceptive, as, despite holding assets that appear to have a significant market value at a given time, their actual worth is effectively zero since they cannot be sold. Thus, while these coins may seem extremely promising for a uninformed investor, they do not reflect a realizable profit in practice. Surprisingly, even with the sellable coins, the strategy shows positive results. Although the negative profit percentage is higher (31.98%), the total profit still reaches 16.43 ETH with a total investment of 35.18 ETH, leading to a percentage profit of 46.69%. Although significantly lower than honeypots, these results show that even a more conservative approach seems to yield profits.
At this point, a spontaneous question arises: are we missing something, or is it really this easy to profit from the creation of new coins, even with a simple strategy, provided that we are (which is not trivial) able to distinguish between honeypots and sellable coins?

### v2

# Price manipulation schemes of new crypto-tokens in decentralized exchanges

@@ Abstract
Blockchain technology has revolutionized financial markets by enabling decentralized exchanges (DEXs) that operate without intermediaries. Uniswap V2, a leading DEX, facilitates the rapid creation and trading of new tokens, which offer high return potential but exposing investors to significant risks. In this work, we analyze the financial impact of newly created tokens, assessing their market dynamics, profitability and liquidity manipulations. Our findings reveal that a significant portion of market liquidity is trapped in honeypots, reducing market efficiency and misleading investors. Applying a simple buy-and-hold strategy, we are able to uncover some major risks associated with investing in newly created tokens, including the widespread presence of rug pulls and sandwich attacks. We extract the optimal sandwich amount, revealing that their proliferation in new tokens stems from higher profitability in low-liquidity pools. Furthermore, we analyze the fundamental differences between token price evolution in swap time and physical time. Using clustering techniques, we highlight these differences and identify typical patterns of honeypot and sellable tokens. Our study provides insights into the risks and financial dynamics of decentralized markets and their challenges for investors.
Keywords:
New Tokens, Uniswap V2, Honeypot, Rug Pulls, Sandwich Attacks, Dynamic Time Warping
@@ 1 Introduction
The advent of blockchain technology has revolutionized financial markets, enabling decentralized platforms that operate independently of centralized intermediaries. Among these innovations, Decentralized Exchanges (DEXs) have emerged as key players in the cryptocurrency ecosystem. Unlike centralized exchanges, DEXs rely on decentralized protocols implemented via smart contracts deployed on blockchains, allowing trustless, peer-to-peer trading without the need for custodial services. This framework has led to the proliferation of cryptocurrency tokens and trading strategies, driven by the innovative mechanism of Automated Market Makers (AMMs)[1], which replaces the traditional order book model with a liquidity pool-based system. One notable example is Uniswap V2 [2, 3], a widely-used DEX protocol that supports the creation of liquidity pools for token pairs, typically involving Wrapped Ethereum (WETH). On average, $\sim 15$ new tokens[4] paired with Ethereum are introduced hourly on Uniswap V2 between October 2, 2024, and December 2, 2024. Many of these tokens initially exhibit extremely high returns, creating an illusion of huge profits and attracting inexperienced investors [5, 6]. However, beneath this apparent simplicity and profitability lie significant risks and vulnerabilities that can lead to partial or total loss of funds [7, 8, 9]. Indeed, the decentralized nature of DEXs introduces a set of security challenges, which have been extensively analyzed in the context of computer security and formal verification methods [10]. These studies highlight that these coins and many investments hide important vulnerabilities, exposing users to risks that require careful evaluation [11]. Among these, the so-called honeypot contracts—first formally defined in [12]—are deceptive smart contracts that mislead users by appearing profitable, encouraging deposits that cannot later be withdrawn due to hidden restrictions, effectively trapping the funds. For instance, a honeypot may include concealed conditions or access controls that appear harmless on interfaces such as Etherscan, but are enforced during execution. These mechanisms often rely on hidden backdoors, such as black- or whitelists, sell taxes, or non-renounced ownership [13, 14].
@@
Contribution of the work. To our knowledge, most existing studies on the vulnerabilities and risks associated with decentralized exchanges, including phenomena such as honeypots and backdoors, have primarily approached the topic from the perspective of computer security and formal verification[15]. Although these works have significantly advanced methods to detect malicious contracts and identify vulnerabilities in smart contracts[16, 17], they often lack a quantitative and financial perspective [18]. Specifically, numerous methodologies have been proposed to identify honeypots or detect vulnerabilities, leveraging static and dynamic analysis techniques, formal methods, and auxiliary service data [19, 20, 21, 22, 12, 13, 14]. However, there has been little to no contribution focused on understanding the broader financial impact of these phenomena, particularly in relation to the scale and significance of the proliferation of new tokens created within the decentralized finance ecosystem [23]. In this work, we aim to bridge this gap by performing a quantitative and financial analysis of the creation and proliferation of new tokens on Uniswap V2.
@@ 2 Dataset and Token Classification
To conduct our analysis, we collected on-chain data from the Ethereum blockchain, focusing on newly created tokens and their trading activity on DEX. The dataset consists of 17,194 tokens launched between October 2, 2024, and December 2, 2024. Data collection was carried out by connecting to an Ethereum node using an RPC provider such as Infura or Alchemy via Web3. Newly created pairs were identified by monitoring blockchain events and tracking the creation of token pairs, recording their contract addresses along with the associated tokens (token_x, token_y). This allowed us to follow the introduction of new tokens into the ecosystem. We then collected corresponding swap events to capture the amounts swapped in and out, providing insights into price dynamics and trading activity. Additionally, we extracted liquidity-related data by examining mint and burn events, recording the amounts of liquidity tokens minted and burned to assess liquidity provisioning over time. This dataset serves as the foundation for analyzing token behavior, price evolution, and liquidity trends of early-stage token dynamics in DEX.
We neglect the new tokens created on Uniswap V3 because only about 1,107 were created within our time window, which is an order of magnitude lower than on Uniswap V2. We also disregard possible interactions with other V3 pools involving the same new tokens created on Uniswap V2, since they amount to approximately 273 tokens-representing a very small fraction. Moreover, most of these tokens are associated with low-liquidity V3 pools.
@@
We performed a time-based analysis, observing each token’s entire trading history to detect any transition into honeypot behavior. We leveraged external security detection services such as Honeypot.is, which specializes in identifying honeypots, and GoPlus Labs, which provides security scoring for tokens. These services allowed us to systematically verify whether a token exhibited selling restrictions at any point in its lifecycle. Based on this analysis, we categorized tokens at the end of their observed lifetime. Thus, we labeled as honeypots those coins that were either initially classified as such or became honeypots later in their lifecycle. However, our analysis indicates that the majority of honeypots transition into this state within the first few swaps. It is important to note that GoPlus and Honeypot.is use different methods to identify honeypots. The former analyzes the smart contract directly, while the latter simulates sell transactions to determine whether a token can be liquidated. GoPlus primarily performs static code inspection and applies heuristic blacklists to detect suspicious or malicious functions within the contract, while Honeypot.is adopts a dynamic approach by simulating real sell transactions to verify whether liquidation is actually possible. The two methods are therefore complementary: static analysis may detect inactive or hidden traps in the code that never trigger on-chain, whereas dynamic simulation can capture behavioral restrictions that cannot be inferred from bytecode alone. In our study, we classify a token as a honeypot if at least one of these two providers flags it as such.
Tokens that never showed honeypot behavior were classified as sellable, comprising approximately 2,080 tokens ($\simeq 12\%$ of the total). In contrast, tokens that showed locked or blocked sales at any stage were classified as honeypots, accounting for around 15,114 tokens ($\simeq 88\%$ of the total).We find that Honeypot.is detects 91.6% of them, while GoPlus Labs detects 81.6%. This shows a substantial agreement between the two providers and shows the robustness of our choice of considering their union.
@@ 3 Early-Stage Properties of Newly Created Tokens
In this section, we analyze the early characteristics of newly created tokens on Uniswap V2, including their creation frequency, classification as sellable tokens or honeypots, and average lifespan. Honeypots significantly distort market dynamics by inflating trading volumes while restricting liquidity. By preventing selling, they create an illusion of high demand and price appreciation, misleading traders and absorbing substantial market volume. Understanding how much liquidity remains locked in these contracts provides insights into the true market potential of sellable tokens and the challenges faced by legitimate projects. A new token lifecycle begins with its creation on the Ethereum blockchain, followed by the establishment of a liquidity pool on Uniswap V2, where the owner pairs the token with WETH to enable trading. Initial trades occur almost immediately [24], often within the first block, driven by automated bots and, in some cases, the creator simulating demand. This early activity frequently causes a sharp price surge due to low liquidity. However, this surge is often short-lived, as subsequent selling pressure, sandwich attacks, and liquidity rug-pulls emerge. Analyzing these early-stage price dynamics, which is precisely what we do in this paper, helps distinguish between tokens with genuine growth potential and those exhibiting market manipulation [25].
@@
(b) WETH amount distribution of all the swaps.
Figure 1: Normalized distributions
@@
Remarkably, the number of tokens created daily is on average around 200, highlighting the sustained activity in token creation (see Appendix A). This trend underscores the importance of Uniswap V2 as a platform to launch and experiment with new tokens. The orange part of the histogram represents the portion of honeypots among these newly created coins. The presence of a substantial fraction of honeypots within this continuous flow of token creation suggests that a significant portion of new projects are designed with malicious intent, aiming to trap liquidity rather than fostering genuine market participation. This dynamic not only skews the perception of overall market activity, but also presents a major challenge for traders and investors, who must navigate an environment where a large share of tokens are inherently unsellable. As a direct consequence, a considerable portion of the liquidity that circulates in the market is effectively locked within honeypot contracts, rendering it inaccessible to traders. This reduces the actual capital available for legitimate trading and investment, further distorting market efficiency. Thus, while the high daily token creation rate may signal strong market engagement, it also reflects an ongoing risk where a substantial amount of liquidity is continuously funneled into fraudulent contracts rather than supporting genuine projects. The peak of the distribution of the tokens number created within one-hour bins occurs around 15. However, the overall distribution is largely dominated by honeypots, which are created more frequently. In contrast, sellable tokens are launched at a significantly lower rate.
@@
A key question is how many swaps newly created tokens experience on average and whether they sustain trading activity or quickly become inactive. This behavior differs between honeypots and sellable tokens. Since honeypots prevent selling, they often see initial high activity as traders unknowingly buy in before realizing that they cannot exit. In contrast, sellable tokens allow for continuous trading, potentially leading to longer lifetimes. Figure 1(a) confirms this: Honeypots exhibit a sharp peak at low swap counts before rapidly decaying, indicating short-lived activity. sellable tokens, while also peaking at low swaps, show a broader distribution, suggesting prolonged trading. This is evident in the median swap count, which is 42 for honeypots but significantly higher at 363 for sellable tokens. Figure 1(b) presents the distribution in swap size in WETH. The sell (negative) side appears uniform, while the buy (positive) side shows distinct peaks at specific WETH values, likely due to bots executing trades with fixed investment amounts. This highlights that a substantial portion of market liquidity remains trapped in honeypots, misleading traders, and distorting the perceived trading volume. It is worth noting that the sell-side transactions observed in honeypots typically originates from privileged wallets under the control of the token creator, who may execute or simulate sales to create the illusion of legitimate market activity.
@@ 4 A buy-and-hold strategy on new tokens
In this section, we investigate the potential profitability of newly created tokens under the assumption that an investor is able to enter a position early in the lifecycle of the token, specifically, at the 60th swap. By systematically applying this strategy across the entire set of newly launched tokens, we can extract valuable insights into their financial behavior. This choice of the 60th swap as entry point is motivated by structural and empirical considerations. Entering earlier, within the first block of transactions, would expose the investor to unrealistically high gas fees and blockchain congestion. Given that the average number of swaps per block at launch is around 27, the 60th swap allows us to avoid the firsts blocks, such that the participation becomes feasible and transaction costs are lower. Moreover, as shown in Figure 1(a), most tokens—especially honeypots—cease trading activity after only a few swaps (median $\sim$42), so starting slightly later allows us to exclude the large fraction of tokens that “die” almost immediately. Finally, as detailed in Appendix B, we confirm empirically that the expected profit function is monotonically increasing when the entry threshold is set at the 60th swap, supporting this value as a reasonable choice. Of course, we acknowledge that the selection of 60 swaps retains a degree of arbitrariness. However, since profits are dominated by extreme honeypot returns, as we shown in this Section, varying the entry threshold within a reasonable range (e.g., 50–100 swaps) does not materially alter the qualitative results. Therefore, the 60-swap threshold can be regarded as a stable and representative benchmark for our analysis.
@@
Our trading bot operates in swap time. In particular, it observes transactions up to the 60th swap before performing its first purchase of tokens using an initial investment of $\Delta x=0.016$ WETH, buying an amount $\Delta y$ of the new token computed using the constant product rule formula
@@
where $x_{60}$ is the reserve of the newly created tokens in the liquidity pool at the 60th transaction, $y_{60}$ is the correspondent reserve of WETH in the liquidity pool and $r=0.997$ is the effective fee factor, accounting for the 0.3% swap fee. The initial investment $\Delta x=0.016\ \mathrm{ETH}$ was chosen based on the empirical distribution of swap volumes at token launch. This value lies close to the mode of the distribution and therefore represents a typical trade size observed on Uniswap V2, ensuring a realistic order magnitude while keeping market impact reasonably small.
After buying $\Delta x$ tokens, the strategy evaluates subsequent swap steps $n$ to determine an optimal time to sell. For each transaction $n>60$, the potential WETH receivable from selling the tokens ($\Delta x_{n}$) is calculated as
@@
where $x_{n}$ and $y_{n}$ are respectively the token reserves of the token and WETH in the pool at transaction $n>60$. The revenue $S_{n}$ at each transaction from the sale is computed as
@@
where $G$ is the gas fee incurred for each transaction that we take fixed to $G=0.002$ WETH. The value G=0.002WETHG=0.002\ $\mathrm${{W}ETH} corresponds to the average cost of a standard Uniswap V2 swap consuming roughly $1.1\text{–}1.5\times 10^{5}$ gas at a price-per-gas of 12–20 gwei, which reflects typical post-launch network conditions.[26, 27] This level implies an average fee of about 4–6 USD per swap (for WETH $\simeq 2,500$ USD), consistent with historical gas costs on Ethereum. Since gas represents a fixed per-transaction cost, its relative impact on profitability rapidly diminishes for sufficiently large trades or returns, becoming negligible in percentage terms. The strategy uses predefined thresholds to decide the swap time $n$ when to sell the tokens:
@@
The same is accounted if the order flow of the token ends before one of these conditions is reached. The choice of these thresholds is informed by the empirical trading lifetimes observed in Figure 1(a). While honeypots typically cease activity after only a few dozen swaps (median $\sim$42), sellable tokens display a median activity of about 363 swaps, and the vast majority stop trading by roughly the 400th swap. It is therefore reasonable to define selling windows within this range, progressively lowering the profit multipliers (2, 1.5, 1.2) as the swap count increases to emulate a decreasing risk appetite over time. These cutoffs are to some extent arbitrary, yet their impact on the aggregate outcomes is limited: given the extremely high returns and the dominance of honeypots that cannot be sold, moderate shifts in the thresholds (e.g., $\pm 50$ swaps or $\pm 0.2$ in the multipliers) do not materially affect the qualitative results.
Figure 2: Distribution of profits in WETH generated by our strategy. Among the 7,213 tokens in which the algorithm invested, 2,101 ($\sim 29.1\%$) failed to exit their position, resulting in a fixed loss of $P_{\text{loss}}=-0.018$ WETH. All remaining tokens yielded a positive profit. To improve visualization, we restrict the distribution to the range between 0 and 0.015 WETH. Beyond this threshold, there are an additional 162 tokens reaching a maximum profit of 1,056.07 WETH.
We apply this strategy to all the coins. The overall combined results indicate that $\sim$ 70.9% of the coins give a positive profit, while approximately $29.1\%$ of the tokens result in the fixed negative profit of $P_{\text{loss}}=-0.018$ WETH. The profit distribution for all tokens is shown in Figure 2, with the range restricted to $P\in[0,0.015]$ to improve visualization. We do not display the negative histogram at $P=P_{\text{loss}}$ and positive outliers. The maximum profit recorded on one coin is 1056.07 WETH. The total profit over all the coins, namely
@@
reaches 2180.24 WETH, with a total investment of 140.06 WETH, leading to an impressive percentage profit of 1556.67%. As expected, however, a large percentage of it is eaten by honeypots.
Indeed, focusing specifically on the honeypot results, we observe an even higher positive percentage of 71.83%, which includes the maximum profit of 1056.07 WETH. The total profit reaches 2163.81 WETH on a total investment of 104.88 WETH, leading to a percentage profit of 2063.13%. This clearly demonstrates that the majority of the apparent profit is generated by honeypot coins, creating the illusion of high profitability. However, this is entirely deceptive, as, despite holding assets that appear to have a significant market value at a given time, their actual worth is effectively zero since they cannot be sold. Thus, while these coins may seem extremely promising for a uninformed investor, they do not reflect a realizable profit in practice. Surprisingly, even with the sellable coins, the strategy shows positive results. Although the negative profit percentage is higher (31.98%), the total profit still reaches 16.43 WETH with a total investment of 35.18 WETH, leading to a percentage profit of 46.69%. Although significantly lower than honeypots, these results show that even a more conservative approach seems to yield profits. We also verified the robustness of our results by varying the initial parameter space — including the entry swap threshold, gas fee assumptions, and exit conditions of the strategy. Despite these changes, the qualitative behavior of the results remains unchanged, owing to the exceptionally large returns observed across the tokens. We also tested an extreme scenario with a gas fee of G = 0.008 ETH, corresponding to a network price of approximately 66 gwei for a 120,000-gas swap—values typical of highly congested blocks right after token deployment. Note that during the period considered in this paper the maximum average gas price in one day has been 52.65 gwei [26, 27] . Even under these extreme conditions, the overall percentage profit remains very large, reaching about 1074%. We also note that small variations of this threshold (e.g., 50 or 100 swaps) do not materially affect the qualitative results, as the average profits shown in Figure 2 remain extremely large. In particular, using the same inputs, the average return is approximately 722% when entering at the 50th swap, 575% at the 70th, and 495% at the 100th. Hence, although the 60th swap provides a theoretically and empirically sound baseline, varying this value leads to very similar conclusions. At this point, a spontaneous question arises: are we missing something, or is it really this easy to profit from the creation of new coins, even with a simple strategy, provided that we are (which is not trivial) able to distinguish between honeypots and sellable coins?